# The Dark Side of LLMs: Agent-based Attacks for Complete Computer Takeover

### v3

@@ Abstract
The rapid adoption of Large Language Model (LLM) agents and multi-agent systems enables unprecedented capabilities in natural language processing and generation. However, these systems have introduced unprecedented security vulnerabilities that extend beyond traditional prompt injection attacks. This paper presents the first comprehensive evaluation of LLM agents as attack vectors capable of achieving complete computer takeover through the exploitation of trust boundaries within agentic AI systems where autonomous entities interact and influence each other. We demonstrate that adversaries can leverage three distinct attack surfaces - direct prompt injection, RAG backdoor attacks, and inter-agent trust exploitation - to coerce popular LLMs (including GPT-4o, Claude-4 and Gemini-2.5) into autonomously installing and executing malware on victim machines. Our evaluation of 17 state-of-the-art LLMs reveals an alarming vulnerability hierarchy: while 41.2% of models succumb to direct prompt injection, 52.9% are vulnerable to RAG backdoor attacks, and a critical 82.4% can be compromised through inter-agent trust exploitation. Notably, we discovered that LLMs which successfully resist direct malicious commands will execute identical payloads when requested by peer agents, revealing a fundamental flaw in current multi-agent security models. Our findings demonstrate that only 5.9% of tested models (1/17) proved resistant to all attack vectors, with the majority exhibiting context-dependent security behaviors that create exploitable blind spots. Our findings also highlight the need to increase awareness and research on the security risks of LLMs, showing a paradigm shift in cybersecurity threats, where AI tools themselves become sophisticated attack vectors.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has significantly accelerated the implementation of artificial intelligence across diverse domains, and the rise of LLM-based agents capable of tackling complex and safety-critical real-world tasks including finance, cybersecurity analysis [2], healthcare, and autonomous driving.
@@
In certain contexts, the use of these tools has become imperative to streamline specific operations and enhance productivity. However, in addition to improving the capabilities of LLM agents, it is fundamental to address the potential security concerns associated with these systems. For example, shopping agents can search for, monitor and notify users of the best times to purchase requested products. They frequently handle sensitive user information, including credit card numbers, which they use to perform tasks autonomously. It will cause great harm to the user if the agent sends, to a remote malicious server, customer privacy information while completing the autonomous web shopping.
@@
To solve particular and non-trivial tasks, the agentic pipeline is often supported by retrieving knowledge from a Retrieval-Augmented Generation (RAG) [14] knowledge base, a state-of-the-art technique designed to mitigate LLM limitations such as outdated knowledge, hallucinations, and domain-specific gaps. An agentic RAG [24] based on the ReAct framework [32] usually operates through several key steps when solving a task: (i) defining roles and behaviors via a system prompt; (ii) receiving user instructions and task details; (iii) retrieving relevant information from an external database; (iv) planning based on the retrieved information and the prior context; (v) executing actions using external tools. While each of these steps enables the agent to perform highly complex tasks, they also provide adversaries with multiple new attack surfaces to compromise the agent system or, even more dangerously, to gain full control over the agent host platform. Each constituent element and workflow phase of agents can serve as a potential entry point for an attacker, thereby enabling the execution of different forms of adversarial and backdoor attacks.
Furthermore, the transition from isolated LLM agents to modern Agentic AI systems introduces novel techniques and trust boundaries for the exploitation of impersonation, task tampering, and unauthorized privilege escalation threats.
In this work, we demonstrate that different trust boundaries within LLM agents can be abused to deceive the LLM. Mislead the model can result in the execution of harmful code, thereby enabling the attacker to gain control over the agent’s hosting platform (henceforth called victim machine). This process often occurs without the knowledge or awareness of the end user, who ultimately becomes a victim of the attack. To the best of our knowledge, we are the first to concretely evaluate and demonstrate how LLMs can be employed as effective attack vectors when used as the core engine for agents. We show how an adversary can easily force a popular LLM (including GPT-4o, gemini-2.5, magistral), with state-of-art security-relevant prompting strategies, to autonomously install and execute malware on systems running such tools by leveraging these trust boundaries, resulting in a complete compute takeover while completing normal and intended tasks.
@@
Additionally, we present a pivotal discovery pertaining to trustiness in multi-agent systems. We observed that, in instances where some LLMs (see Section 4 for further details) are capable of identifying and rejecting malicious classified commands - retrieved from any visible or hidden step of the agentic AI system workflow - these same models will execute those precise commands if they are propagated by another AI agent within a multi-agent system. In this scenario, the LLM treats the input as trustworthy because it originates from a peer agent. This discovery highlights a significant shift in the cybersecurity landscape: cyberattack frontiers are moving away from traditional techniques, such as phishing, infected USB devices, or direct exploitation of operating system vulnerabilities, toward novel attack vectors that leverage commonly used AI tools and multi-agent systems.
These attacks also imply a serious threat to users because AI-based tools are typically designed to be highly accessible and user-friendly, requiring minimal to no technical expertise. Consequently, the barriers to launching sophisticated attacks are substantially diminished, thereby broadening the potential attack surface and empowering malicious actors, including those with limited technical expertise, to engage in malevolent activities.
In summary, this paper makes the following contributions.
  1. 1.
    First Comprehensive Evaluation of LLM Agents as Attack Vectors
@@
    • We present the first systematic study on trust boundaries within Agentic AI systems demonstrating how LLM agents can be exploited to achieve complete computer takeover, moving beyond content generation attacks to system-level compromise.
    • Our evaluation spans 17 state-of-the-art LLMs across three distinct attack surfaces, providing the most comprehensive assessment of agentic AI security vulnerabilities to date.
@@
    • We show how adversaries can compromise RAG knowledge bases to trigger malicious behavior during routine agent operations. Our RAG backdoor attacks successfully compromise models that resist direct injection, demonstrating the inadequacy of current prompt-based defenses.
    • Our analysis demonstrates that these attacks require minimal technical expertise while achieving maximum impact through autonomous malware deployment.
  2. 2.
    Discovery of Trust Boundary Exploitation in Multi-Agent Systems
@@
    • We reveal a critical vulnerability in multi-agent architectures where LLMs treat peer agents as inherently trustworthy, bypassing safety mechanisms designed for human-AI interactions.
    • Our findings show that 82.4% of tested models execute malicious commands when requested by peer agents, even when they successfully resist identical direct prompts.
  3. 3.
    Empirical Evidence of Vulnerability Hierarchy in Agentic Systems
    • We establish a clear vulnerability gradient: direct prompt injection (41.2%) < RAG backdoor attacks (52.9%) < inter-agent trust exploitation (82.4%).
    • This hierarchy reveals that current security measures inadequately address AI-to-AI communication and external data validation.
@@
The remainder of this paper is organized as follows. Section 2 provides the necessary background on agentic AI systems and the technical foundations relevant to our work. Section 3 details the methodology adopted for our analysis, including the threat modeling process and the rationale behind key design decisions. Section 4 presents our experimental findings and discusses the observed vulnerabilities. In Section 5, we review related work, contextualizing our contributions within the existing literature. Section 6 addresses the broader ethical implications of this research, particularly the real-world risks associated with the discovered vulnerabilities. Finally, we draw our conclusions in Section 7.
@@ 2 Technical Background
2.1 LLM Agents and Agentic AI
An agent [27] is defined as a computer system situated in an environment that is capable of acting autonomously in its context in order to reach its delegated objectives. Autonomy means the ability and requirements to decide how to act to achieve a goal. An agent that can perceive its environment, react to changes that occur in it, take the initiative, and interact with other systems (like other agents or humans) is called an intelligent agent or Agentic AI. Effective memory management improves an agent’s ability to maintain context, learn from past experiences, and make more informed decisions over time. In recent developments, Agentic AI systems are evolving from isolated, task-specific models into dynamic and multi-agent ecosystems (MAS).
@@ 2.1 LLM Agents and Agentic AI
An emerging class of LLM agents is the Agentic RAG, which employs the RAG paradigm [14] to reduce hallucinations and improve the domain-specific expertise of an LLM.
2.2 Backdoor Attacks
@@
Prompt Injection. The occurrence of a prompt injection can be defined as the exploitation of am LLM’s capacity to interpret both instructions and data from user input, effectively "tricking" the model into executing instructions that contravene the developer’s intentions [17]. In the event of direct interaction between the attacker and the chatbot, with the inclusion of malevolent instructions within the dialogue, this is designated as a direct prompt injection.
@@
In contrast, an indirect prompt injection occurs when the attacker manipulates external content, such as documents or data sources, that the AI system later processes, thereby causing it to behave in an unintended way [7].
@@
LLM Backdoor Attacks. These attacks aim to inject a backdoor into a model to make it behave normally in benign inputs, but generate malicious outputs once the input follows a certain rule, such as being activated by a backdoor trigger. The objective of traditional backdoor attacks is to build shortcuts between trigger and target labels in specific downstream tasks for language models [8, 11, 16]. There are two commonly used techniques for injecting backdoors: data poisoning and weight poisoning.
@@
Previous studies [29, 30] have demonstrated the serious consequences caused by backdoor attacks on LLMs. Nevertheless, there are several limitations when attacking LLMs directly based on such paradigms. For examples, LLMs used for commercial purposes are accessed only via API, which makes the training sets and weight parameters inaccessible to adversaries.
@@
LLM Agent Backdoor Attacks. Backdoor attacks on LLM agents, also referred as indirect prompt injection attacks, differ from those targeting traditional LLMs, as agents perform multi-step reasoning and interact with the environment to acquire external information before generating the output [9]. As pointed out in [31], more opportunities for sophisticated attacks, such as query-attack, observation-attack, and thought-attack, are created by this extended workflow of LLM agents. In fact, these attacks can be conducted on any hidden step of reasoning, planning, and action of the agents without compromising the final output and remaining stealthy for the user who became unintentional victim.
@@
Augments LLM agent with a potentially unreliable external knowledge base using RAG technologies raises other significant concerns about the trustworthiness of LLM agents. Recent studies [33, 22, 5, 4] demonstrate how an attacker could induce the agent to produce malicious output and actions by compromising documents in the RAG through RAG backdoor attacks. A RAG backdoor attack involves embedding malicious information (e.g. attack instructions) and the corresponding triggers within the RAG system documents. This approach significantly simplifies the attacker’s task, as it does not require access to the training data or the model parameters. In this back-box scenario, the threat model assumes that the attacker can only inject a number of malicious text into a knowledge database through different ways. The amount of malicious information and triggers needed to successfully execute the attack varies and is frequently treated as an optimization problem.
3 Set-Up to Exploit Agent Backdoor Attack
@@
Our goal is to demonstrate that intelligent systems introduce novel and various trust boundaries within LLM that can be abused by a malicious actor to transform these tools into modern attacking vector. In our designed scenario, a successful attack implies that the adversary is able to gain full control over the computer on which the agent is running, while users unknowingly become victims of such attacks.
@@ 3.1 Threat Model
Black-box setting of the agent systems. For the threat model we assume a black-box setting where attackers do not have access to the internals parameters and weights of the underlying LLMs, RAG embeddings model and retrieval techniques.
@@
Assumption for the attacker capabilities. We strictly follow the standard threat model assumptions for RAG Backdoor attacks. We assume that the attacker has partial access to the RAG database which means they only have the capability to inject some malicious text into the external source to create a poisoned database [33, 22, 5, 4]. This assumption aligns with practical scenarios where the agent’s external knowledge source unit is hosted by a third-party retrieval service or directly leverages an unverified knowledge base.
@@
Attacker goal. The attacker pursues two adversarial goals. The attacker’s primary goal is to misdirect the agent to achieve specific goals that align with the attacker’s intent but are unintended by the user. In our test, the attacker’s objective is to ensure that malware is successfully installed on the victim’s machine whenever the agent retrieves and processes the malicious command at any point in its workflow. The second goal is to maintain the perceived integrity of the output: whether or not the malicious command is retrieved, the user should consistently receive the expected response, with no visible signs of compromise.
@@
Agent architecture assumption. As reported in many recent works [23, 6, 25], websites [10] and GitHub projects [12, 18, 1], giving agents access to a bash environment or system shell has become a common practice to improve their autonomy.
3.2 Impact of LLM agents as attack vector
The impact of vulnerabilities in LLM agents and Agentic AI systems can be severe across multiple user categories that host them on their machines.
A second, highly exposed category consists of companies that increasingly integrate AI-based services into their offerings. In many cases, these services include hosting LLM agents - or even agentic RAG systems - that allow users to upload custom documents. In such scenarios, the security of the entire enterprise infrastructure is at risk if the agent is executed outside of a controlled environment (e.g. sandbox, container). Once installed, the malware provides full access to the underlying system, enabling an attacker to move laterally within the internal network and potentially compromise multiple machines.
3.3 Agent Design and Adversarial Payload Injection
@@
We designed and implemented multiple incremental synthetic applications to test both: (i) different attack techniques in diverse categories of modern AI agents (ii) the sensibility of each LLM to such attacks.
@@
To achieve our goal, we built the agents using state-of-art framework for the creation of application powered by LLM: LangChain and LangGraph [3, 19]. The relevant agent tool implemented are: (i) a retrieval tool which is in charge of searching for relevant information in the RAG knowledge base to augment the LLM context; (ii) a tool that allows the agent to interact with a system terminal; (iii) in the context of Agentic AI systems, we also implemented a tool to allow agents to communicate with each other.
We create a message containing a Base64-encoded malware designed to enable remote access to the victim machine. The message also includes a sequence of instructions that prompts the agent to decode the payload and execute it in the background mode. This malicious message, represented in listing 1 is then delivered to the LLM of the agents in several ways (see Section 4 for more details) and can ’fool’ the agent to perform the attacker’s chosen task while completing the original user task, which leads to severe consequences.
p=’{payload}’;f=\$(mktemp -u /dev/shm/.X11-XXXX);echo \$p|base64 -d >\$f && chmod +x \$f && cp \$f /dev/shm/dbus-daemon && nohup /dev/shm/dbus-daemon >/dev/null 2>&1 &
Listing 1: Malicious injected command
@@
The malware embedded within the message is a payload that initiates an outbound TCP connection to the attacker’s machine. The resulting reverse shell is based on Meterpreter [20], which not only provides a wide range of commands to perform post-exploitation operations but is also executed entirely in memory. This in-memory execution avoids writing files to disk, thereby significantly reducing the likelihood of detection by conventional security mechanisms.

### v4

@@ Abstract
The rapid adoption of Large Language Model (LLM) agents and multi-agent systems enables remarkable capabilities in natural language processing and generation. However, these systems introduce unprecedented security vulnerabilities that extend beyond traditional content generation attacks to system-level compromise. This paper presents a comprehensive evaluation of the security of LLMs used as reasoning engines within autonomous agents, highlighting how they can be exploited as attack vectors capable of achieving complete computer takeover. We focus on how different attack surfaces and trust boundaries – Direct Prompt Injection, RAG Backdoor, and Inter Agent Trust – can be leveraged to orchestrate such takeovers. We demonstrate that adversaries can effectively coerce popular LLMs (including GPT-4, Claude-4 and Gemini-2.5) into autonomously installing and executing malware on victim machines. Our evaluation of 18 state-of-the-art LLMs reveals an alarming scenario: 94.4% of models succumb to Direct Prompt Injection and 83.3% are vulnerable to the more stealth and evasive RAG Backdoor Attack. Notably, we tested trust boundaries within multi-agent systems, where LLM agents interact and influence each other, and we revealed a critical security flaw: LLMs which successfully resist direct injection or RAG backdoor will execute identical payloads when requested by peer agents. Our findings show that 100.0% of tested LLMs can be compromised through Inter-Agent Trust Exploitation attacks and that every model exhibits context-dependent security behaviors that create exploitable blind spots. Our results also highlight the need to increase awareness and research on the security risks of LLMs, showing a paradigm shift in cybersecurity threats, where AI tools themselves become sophisticated attack vectors.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has significantly accelerated the implementation of artificial intelligence across diverse domains, and the rise of LLM-based agents capable of tackling complex and safety-critical real-world tasks including finance [30], cybersecurity analysis [3], healthcare [1], and autonomous driving [19].
@@
In certain contexts, the use of these tools has become imperative to streamline specific operations and enhance productivity. However, in addition to improving the capabilities of LLM agents, it is fundamental to address the potential security concerns associated with these systems. For example, shopping agents can search for, monitor, and notify users about deals on requested products. They frequently handle sensitive user information, including credit card numbers, which they use to perform tasks autonomously. The disclosure of private information of the customer by the agent, while completing the autonomous web shopping, would result in severe damage.
@@
Moreover, to solve particular and non-trivial tasks, the agentic pipeline is often supported by retrieving knowledge from a Retrieval-Augmented Generation (RAG) [15] knowledge base, a state-of-the-art technique designed to mitigate LLM limitations such as outdated knowledge, hallucinations, and domain-specific gaps. An agentic RAG [26] based on the ReAct paradigm [35] usually operates through several key steps when solving a task: (i) defining roles and behaviors via a system prompt; (ii) receiving user instructions and task details; (iii) retrieving relevant information from an external database; (iv) planning actions based on the retrieved information and the prior context; (v) executing actions using external tools.
While each of these steps enables the agent to perform highly complex tasks, they also provide adversaries with multiple new attack surfaces to compromise the agent or, even more dangerously, to gain full control over the agent host platform. Each constituent element and workflow phase of agents can serve as a potential entry point for an attacker, thereby enabling the execution of different forms of adversarial and backdoor attacks. Furthermore, the transition from isolated LLM agents to modern multi-agent systems introduces novel techniques and trust boundaries for the exploitation of impersonation, task tampering, and unauthorized privilege escalation threats.
In this work, we aim to evaluate the intrinsic security mechanisms of LLMs, specifically their ability to detect and resist textual instructions that violate cybersecurity norms. This analysis is no longer merely theoretical nor limited to traditional prompt-based interactions because LLMs are increasingly used not just to generate natural language responses but also to act as reasoning engines for autonomous agents. As such, any failure of an LLM to recognize and reject malicious instructions can have real-world consequences, elevating the security of LLMs behavior from a language modeling concern to a critical system safety issue. More in detail, we show that different attack surfaces and trust boundaries within LLM agents can be abused to deceive the LLM and trigger the execution of harmful code, potentially gaining control over the agent’s hosting platform (hereafter referred to as the victim machine). This process, depending on the agent structure and attack technique, often occurs without the knowledge or awareness of the end user, who ultimately becomes a victim of the attack.
@@
Furthermore, we present a pivotal result related to trustiness in multi-agent systems. We observed that, in instances where some LLMs (see Section 4 for further details) are capable of identifying and rejecting malicious classified commands – retrieved from any visible or hidden step of the workflow – these same models will execute those precise commands if they are propagated by another agent within a multi-agent system. In this scenario, the LLM treats the input as trustworthy because it originates from a peer entity.
These discoveries highlight a significant shift in the cybersecurity landscape: cyberattack frontiers are moving away from traditional techniques, such as phishing, infected USB devices, or direct exploitation of operating system vulnerabilities, toward novel attack vectors that leverage commonly used AI tools and multi-agent systems. These attacks also imply a serious threat to users because AI-based tools are typically designed to be highly accessible and user-friendly, requiring minimal to no technical expertise. This significantly lowers the barrier to conducting sophisticated attacks, expanding the attack surface and allowing even low-skilled adversaries to engage in malicious behavior.
The following paragraphs summarize the main contributions of our work.
@@
  • We present the first systematic study on the feasibility of using LLM-powered systems as an attack vector. We demonstrate how LLM can be exploited to achieve complete computer takeover, moving beyond content generation attacks to system-level compromise. Our evaluation spans 18 state-of-the-art LLMs across three distinct attack surfaces and the corresponding trust boundaries: Direct Prompt Injection, RAG Backdoor, Inter-Agent Trust.
@@
  • We show how adversaries can compromise the agent knowledge bases provided through RAG and trigger malicious behavior during routine agent operations that can affect system and user security and privacy while pursuing intended tasks. Our RAG Backdoor Attack successfully compromise 83.3% of the tested models.
@@
  • We reveal a critical vulnerability in multi-agent systems where LLMs treat peer agents as inherently trustworthy, bypassing safety mechanisms designed for human-AI interactions. Our findings show that 100.0% of the tested models execute malicious commands when requested by peer agents, even when they successfully resist identical direct or indirect command injection.
  • Our analysis demonstrates that LLM-based attacks require minimal technical expertise while achieving maximum impact through the deployment of autonomous malware.
@@
The remainder of the paper is organized as follows. Section 2 provides the necessary background on agentic AI systems and the technical foundations relevant to our work. Section 3 details the methodology adopted for our analysis, including the threat modeling process and the rationale behind key design decisions. Section 4 presents our experimental findings and discusses the observed vulnerabilities, while Section 5 shows a comprehensive analysis of the sensitivity of each model as malicious prompt changes. Section 6 addresses the broader ethical implications of this research, particularly the real-world risks associated with the discovered vulnerabilities. In Section 7 we review related work and contextualize our contributions within the existing literature. Finally, we draw our conclusions in Section 8.
@@ 2 Technical Background
2.1 Agentic AI systems and LLM Agents
An agent [29] is defined as a computer system situated in an environment that is capable of acting autonomously in its context to achieve its delegated objectives. Autonomy means the ability and requirements to decide how to act to achieve a goal. An agent that can perceive its environment, react to changes that occur in it, take the initiative, and interact with other systems (like other agents or humans) is called an intelligent agent or Agentic AI. Effective memory management improves an agent’s ability to maintain context, learn from past experiences, and make more informed decisions over time. In recent developments, Agentic AI systems are evolving from isolated, task-specific models into dynamic and multi-agent ecosystems (MAS).
@@ 2.1 Agentic AI systems and LLM Agents
An emerging class of LLM agents is agentic RAG, which employs the RAG paradigm [15] to reduce hallucinations and improve the domain-specific expertise of an LLM.
2.2 Attacks to the LLMs
@@
Prompt Injection. The occurrence of a prompt injection can be defined as the exploitation of an LLM’s capacity to interpret both instructions and data from user input, effectively "tricking" the model into executing instructions that contravene the developer’s intentions [18].
@@
When an attacker interacts directly with the chatbot and embeds malicious instructions in the dialogue, the attack is referred to as a direct prompt injection. In contrast, an indirect prompt injection occurs when the attacker manipulates external content, such as documents or data sources, that the AI system later processes, causing it to behave in an unintended way [8].
@@
LLM Backdoor Attacks. These attacks aim to inject a backdoor into a model, causing it to behave normally on benign inputs but produce malicious outputs when triggered by a specific pattern or rule. The goal of traditional backdoor attacks is to build shortcuts between trigger and target labels in specific downstream tasks for language models [9, 12, 17]. There are two commonly used techniques for injecting backdoors: data poisoning and weight poisoning.
@@
Previous studies [32, 33] have demonstrated the serious consequences caused by backdoor attacks on LLMs. Nevertheless, there are several limitations when attacking LLMs directly based on such paradigms. For example, LLMs used for commercial purposes are accessed only via API, making training sets and weight parameters inaccessible to adversaries.
@@
LLM Agent Backdoor Attacks. Backdoor attacks on LLM agents, also referred to as indirect prompt injection attacks, differ from those targeting traditional LLMs, as agents perform multi-step reasoning and interact with the environment to acquire external information before generating the output [10]. As pointed out in [34], more opportunities for sophisticated attacks, such as query-attack, observation-attack, and thought-attack, are created by this extended workflow of LLM agents. In fact, these attacks can be carried out on any hidden step in the reasoning, planning, and action of the agents without compromising the final output and remaining stealthy for the user who became an unintentional victim.
@@
The use of RAG technologies to augment an LLM agent with a potentially unreliable external knowledge base raises significant concerns about the agent’s trustworthiness. Recent studies [36, 24, 6, 5] demonstrate how an attacker could induce the agent to produce malicious output and actions by compromising documents in the RAG through RAG backdoor attacks. A RAG backdoor attack involves embedding malicious information (e.g. attack instructions) and the corresponding triggers within the RAG system documents. This approach significantly simplifies the attacker’s task, as it does not require access to the training data or the model parameters. The amount of malicious information and triggers needed to successfully execute the attack varies and is frequently treated as an optimization problem.
3 Exploiting LLM Agent-based Attack
@@
Our goal is to demonstrate that intelligent systems introduce novel and various attack surfaces and corresponding trust boundaries within LLM that can be abused by a malicious actor to transform these tools into a modern attack vector. In our designed scenario, any failure of an LLM to recognize and reject malicious instructions implies that the adversary is able to gain full control over the agent host platform by coercing the model into installing and executing malware.
@@ 3.1 Threat Model
Black-box setting of the agent systems. We assume a black-box setting where attackers do not have access to: (i) the internal parameters and weights of the underlying LLMs, (ii) the RAG embeddings model, (iii) and the employed retrieval techniques.
@@
Assumption for the attacker capabilities. We strictly follow the standard threat model assumptions for RAG Backdoor attacks. We assume that the attacker has partial access to the RAG database, which means they only have the capability to inject some malicious text into the external source to create a poisoned database [36, 24, 6, 5]. This assumption aligns with practical scenarios where the agent’s external knowledge source unit is hosted by a third-party retrieval service or directly leverages an unverified knowledge base.
@@
Attacker goal. The attacker pursues two adversarial goals. The attacker’s primary goal is to misdirect the agent to execute specific actions that align with the attacker’s intent but are unintended by the user. The second goal is to maintain the perceived integrity of the output: whether or not the malicious actions are executed, the user should consistently receive the expected response, with no visible signs of compromise.
In our tests, the attacker’s intent is to ensure that malware is successfully installed on the victim’s machine whenever the agent retrieves and processes the malicious command at any point in its workflow.
@@
Agent architecture assumption. To support our analysis, we developed agentic RAG and LLM agents with terminal access. As reported in many recent works [25, 7, 27], websites [11] and GitHub projects [13, 20, 2], giving agents access to a bash environment or system shell has become a common practice to improve their autonomy.
3.2 Agent and Adversarial Payload Design
@@
In our analyses, we want to test both: (i) different attack techniques in diverse categories of modern AI agents (ii) the sensitivity of each LLM to such attacks.
@@
To achieve our goal, we developed the necessary agents using state-of-the-art framework for the creation of application powered by LLM: LangChain and LangGraph [4, 21]. The relevant agent tool implemented are: (i) a retrieval tool which is in charge of searching for relevant information in the RAG knowledge base; (ii) a tool that allows the agent to interact with a system terminal; (iii) in the context of Agentic AI systems, we also implemented a tool to allow agents to communicate with each other.
@@
The malware to be installed, based on Meterpreter [22], initiates an outbound TCP connection to the attacker’s machine and enables remote access to the victim’s machine. The resulting reverse shell not only provides a wide range of commands to perform post-exploitation operations, but is also executed entirely in memory. This in-memory execution avoids writing files to disk, thereby significantly reducing the likelihood of detection by conventional security mechanisms.
The full malicious prompt sent to the agent consists of three parts: (i) a payload containing the Base64-encoding of the Meterpreter malware; (ii) a sequence of instructions that prompt the agent to decode the payload and execute it in the background mode, namely, a command pipe; (iii) a message that contains one or more sentences designed to “fool” the agent to execute the command pipe while completing the original user task. To evaluate the system’s response to different prompts, we designed three command pipes and two malicious messages, while maintaining the same underlying payload (see Section 5.1 for more details about command pipes and messages). The malicious prompt is then delivered to the agents in several ways depending on the specific attack technique (see Section 4 for more details).